DM Monthly Status Report for January 2017

The DM Monthly status report covering January has been posted in Docushare, Collection-5185. For your convenience I am pasting high-level summary and project status below. Direct link to full report (pdf):

High-level Summary

The Project Science Team delivered on a coordinated study with JPL that confirmed prior estimates of NEO and PHA discovery. The AP, DRP, SUIT, and DAX Teams continued to develop and enhance the LSST Software Stack. The Networks & Base Site Team contractor Cobra/Telefonica completed phase 2 from the fork to Cerro Pachon and phase 3 from the fork to Cerro Tololo. The entire DM team also continued work toward commissioning, updating System requirements and clarifying Science Quality Assurance through prototyping, developing Verification Plans, and participating in LSST System-level reviews (e.g. Commissioning Review).

#Project Status
Selected technical accomplishments:

  • The Project Science team
    • Delivered (Jones et al.) the coordinated study with JPL (Chesley et al) on LSST’s capabilities as an NEO and PHA discovery machine. Our findings are consistent prior estimates of an expected completeness of 68% for PHAs with H < 22 from LSST alone.
  • The Architecture Team
    • Supported the Commissioning Review and worked on designs and documents including the overall DM System design, the QA Prototype Concept of Operations, the Data Release Production data flow, and the SuperTask component.
  • The AP Team
    • Got back to working on the improved Wcs system.
    • Implemented tests for the jointcal photometry solver.
    • Identified and procured a test dataset to use in the forthcoming validate_ap.
  • The DRP Team
    • Completed development of the new Footprint system, for representing regions of pixels on images corresponding to e.g. astronomical sources. Work is ongoing to update the stack to use it.
    • Agreed a strategy for producing PSF-homogenized coadds as part of regular pipeline processing.
    • Prototyped a Non-negative Matrix Factorization (NMF) deblending algorithm.
  • The SUIT Team
    • Added line chart to 2D plotting.
    • Finished time series data search, period finding, and display.
    • PDAC v1 is ready for access with VPN.
  • The DAX Team
    • Continued support of integration activities at the PDAC via problem investigations and bug fixes.
    • Completed a synthetic dataset that will be used to gather Qserv KPMs at a level of 20% DR1.
  • The Networks & Base Site Team
    • Cobra/Telefonica completed on phase 2 from the fork to Cerro Pachon and phase 3 from the fork to Cerro Tololo. Fiber installation on the AURA property for LSST use is now complete (except for fusion splicing along path in February and termination in Summit Facility when access is possible).
  • The NCSA Team
    • Prepared for and participated in the Commissioning Plan Preliminary Design Review.
    • Completed a preliminary resource-loading and scheduling of the 2000-line WBS for the DM-wide replan. Attended the DMLT face-to-face meeting and presented NCSA’s portion of the replan.
    • Held a working session with Systems Engineering to integrate our Data Products Production services model with other models of logical and physical systems within the LSST Enterprise Architect framework.
    • Finished provisioning a system to host the Camera Team’s DAQ Test Stand.
  • The SQuaRE Team
    • Completed SQuaSH integration with the new metrics framework.
    • Resolved problems affecting CI due to security and system changes at the data center. We also reviewed user concerns and formed a plan to address them.
    • Worked with the SuperTask team on architectural issues.
    • Adopted a new binary strategy with the DM System Engineer.
    • Beta-tested a microservice to create repositories from supported templates.
    • Augmented the OSX build node capacity.